OpenStudy (anonymous):

.One location of The Smart Growing Preschool has 22 children in the primary class and 18 children in the toddler class. The other location has 30 children in the primary class and 15 children in the toddler class. The revenue at the first location is $24,250, and the revenue at the second location is $27,150. How much do parents pay for a primary child and a toddler child every month? A.$595 for a primary child and $620 for a toddler child. B.989 for a primary child and $1,558 for a toddler child. C.1,353 for a primary child and $307 for a toddler child.

OpenStudy (cj49):

30+15=45 30/45x27150=18100 18100/30=603 for a primary child its 595(the closest option to 603)